OpenStudy (anonymous):

WILL GIVE MEDAL! Which equation is related to the equation shown? 5x = 35 A.x = 35 + 5 B.x = 35 ÷ 5 C.x = 35 − 5 D.x = 35 • 5

OpenStudy (gorv):

divide both side by 5 to find x first

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the answer B?

OpenStudy (gorv):

yeah

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

I have another one

OpenStudy (anonymous):

Which equation represents the word sentence? The product of 4 and a number increased by 6 is 30.

OpenStudy (gorv):

letthe number is x

OpenStudy (gorv):

now increase it by 6

OpenStudy (anonymous):

A.4x + 6 = 30 B.4x – 6 = 30 C4x = 30 + 6 D.4x + 30 = 6

OpenStudy (gorv):

increase x by 6=????

OpenStudy (anonymous):

is it A??

OpenStudy (gorv):

yeah u r right

OpenStudy (anonymous):

one more

OpenStudy (gorv):

post new Q

OpenStudy (anonymous):

ok